# Large Language Models Engineer Too Many Simple Features for Tabular Data

### v1

@@ Abstract
Tabular machine learning problems often require time-consuming and labor-intensive feature engineering. Recent efforts have focused on using large language models (LLMs) to capitalize on their potential domain knowledge. At the same time, researchers have observed ethically concerning negative biases in other LLM-related use cases, such as text generation. These developments motivated us to investigate whether LLMs exhibit a bias that negatively impacts the performance of feature engineering. While not ethically concerning, such a bias could hinder practitioners from fully utilizing LLMs for automated data science. Therefore, we propose a method to detect potential biases by detecting anomalies in the frequency of operators (e.g., adding two features) suggested by LLMs when engineering new features. Our experiments evaluate the bias of four LLMs, two big frontier and two small open-source models, across 27 tabular datasets. Our results indicate that LLMs are biased toward simple operators, such as addition, and can fail to utilize more complex operators, such as grouping followed by aggregations. Furthermore, the bias can negatively impact the predictive performance when using LLM-generated features. Our results call for mitigating bias when using LLMs for feature engineering.
1 Introduction
@@
Machine learning problems for tabular data exist in many domains, such as medical diagnosis, cybersecurity, and fraud detection (Borisov et al., 2022a; van Breugel & van der Schaar, 2024). The original data for these problems (e.g., an Excel sheet) often requires manual feature engineering by a domain expert to solve the machine learning problem accurately (Tschalzev et al., 2024). During (automated) feature engineering, various operators (e.g., Add, Divide, GroupByThenMean) are applied to existing features to create new features (Kanter & Veeramachaneni, 2015; Prado & Digiampietri, 2020; Mumuni & Mumuni, 2024).
@@
Large language models (LLMs) understand various domains (Kaddour et al., 2023; Kasneci et al., 2023; Hadi et al., 2024), tabular data (Ruan et al., 2024; Fang et al., 2024), and feature engineering (Hollmann et al., 2024; Jeong et al., 2024; Malberg et al., 2024). Thus, data scientists have started to leverage LLMs for feature engineering [^1] For example, see these recent Kaggle competition write-ups (Hatch, 2024; Türkmen, 2024).; liberating practitioners from extensive manual labor.
@@
Despite their utility, LLMs are known to have negative biases as observed for chat applications (Kotek et al., 2023; Navigli et al., 2023; Gallegos et al., 2024; Bang et al., 2024) or when "meticulously delving" into text generation (Liang et al., 2024). Observing such biases motivated us to determine whether LLMs also exhibit a bias that negatively impacts the quality of their engineered features. If a bias is found and can be circumvented, LLMs would become a more potent tool for data scientists.
@@
To determine whether a bias exists, we inspect the frequency of operators LLMs use when engineering features. This parallels inspecting the frequency of words to detect LLM-generated text (Liang et al., 2024). Assuming LLMs use their world knowledge and reasoning capabilities to employ the most appropriate operator, we expect the operators’ frequencies to be similar to those of the optimal operators. That is, if adding two features is often the optimal feature, then an LLM would frequently add two features. Moreover, if the LLM does not know the optimal operator, it should resort to a random search over operators.
@@
Therefore, we compare the frequencies of operators used by an LLM to those obtained by searching for the optimal features with automatic black-box feature engineering using OpenFE (Zhang et al., 2023). Using this approach, we evaluated the bias of 4 LLMs, namely GPT-4o-mini (OpenAI, 2024), Gemini-1.5-flash (Gemini-Team, 2024), Llama3.1-8B (Touvron et al., 2023), and Mistral7B-v0.3 (Jiang et al., 2023). We obtained the distribution over operators for 27 tabular classification datasets unknown to all LLMs.
@@
Our results demonstrate that LLMs can have a negative bias when engineering features for tabular data. LLMs favor simple operators during feature engineering (e.g., Add), while some LLMs rarely use more complex operators (e.g., GroupByThenMean). In contrast, automatic black-box feature engineering favors complex operators but also uses simple operators.
@@
In particular, we observed a strong bias and negative impact for GPT-4o-mini and Gemini-1.5-flash, two big frontier models (Chiang et al., 2024). Both select simple operators most often and their generated features decrease the average predictive performance. In contrast, Llama3.1-8B and Mistral7B-v0.3, the small open-source models, are less biased or negatively impacted. Nevertheless, no LLM is close to the distribution over operators obtained by OpenFE. Likewise, the features generated by OpenFE improve the predictive performance on average the most.
@@
2 Related Work
Feature Engineering Without Large Language Models. Previous work dedicated considerable effort toward automating the process of feature engineering (Kanter & Veeramachaneni, 2015; Prado & Digiampietri, 2020; Mumuni & Mumuni, 2024). Various black-box methods have been proposed, such as ExploreKit (Katz et al., 2016), AutoFeat (Horn et al., 2020), BioAutoML (Bonidia et al., 2022), FETCH (Li et al., 2023), and OpenFE (Zhang et al., 2023). These methods typically generate new features in two steps: 1) create a large set of candidate features by applying mathematical (e.g., Add) or functional (e.g., GroupByThenMean) operators to features, and 2) return a small set of promising features selected from all candidate features.
Feature Engineering with Large Language Models. LLMs allow us to exploit their (potential) domain knowledge for feature engineering. LLMs can act as a proxy to a domain expert or data scientist during the feature engineering process. To illustrate, LLMs can be prompted to suggest code for generating new features (Hollmann et al., 2024; Hirose et al., 2024), to select predictive features (Jeong et al., 2024), or to use rule-based reasoning for generation (Nam et al., 2024).
@@ 2 Related Work
Other Applications of Large Language Models for Tabular Data. Many researchers recently started using LLMs for applications related to tabular data. To avoid confusion in this plethora of recent work, we highlight similar but not directly related work to our contribution. Our work is not directly related to LLMs to tabular question answering (Ghosh et al., 2024; Grijalba et al., 2024; Wu et al., 2024), tabular dataset generation (Borisov et al., 2022b; van Breugel et al., 2024; Panagiotou et al., 2024), tabular data manipulation (Zhang et al., 2024; Qian et al., 2024; Lu et al., 2024), or tabular few-shot predictions (Hegselmann et al., 2023; Han et al., 2024; Gardner et al., 2024)
3 Method: Analyzing Feature Engineering Bias
Figure 1: Our Method to Analyze Feature Engineering Bias of an LLM. Our method is split into three stages. In the first stage A), we discard tabular datasets, which the LLM might have memorized. In the second stage B), we instruct the LLM to select the best operators to engineer new features. At the same stage, we use OpenFE (Zhang et al., 2023), black-box automated feature engineering, to determine a proxy for the optimal operators. In the third stage C), we compare the frequencies of operators for the LLM and OpenFE. That is, we look for anomalies when contrasting the distributions, such as using simple operators much more than complex ones.
@@ 3 Method: Analyzing Feature Engineering Bias
The LLM is then instructed to generate precisely one new feature by selecting one or two existing features and applying one of the available operators. We employ chain-of-thought (CoT) prompting (Wei et al., 2023) to boost the expressive power of the LLM (Li et al., 2024). In addition, the LLM explains why a feature was generated with CoT. We also employ a feedback loop, similar to CAAFE (Hollmann et al., 2024), which we detail in Appendix A.
@@
Fundamentally, we aim to compare the distribution over operators suggested by the LLM to the distribution over the optimal operators. That said, we do not know the optimal operators for a dataset. Thus, as a proxy, we use the operators suggested by OpenFE (Zhang et al., 2023).
@@ 4 Experiments
Large Language Models. We use four LLMs hosted by external providers via APIs. In detail, we used GPT-4o-mini (OpenAI, 2024) and Gemini-1.5-flash (Gemini-Team, 2024) to represent big frontier LLMs and Llama3.1 8B (Touvron et al., 2023) and Mistral 7B Instruct v0.3 (Jiang et al., 2023), hosted by Together AI [^3] https://www.together.ai/, to represent small open-source models. The API usage costed ${\sim}200\$$.
@@
Datasets. We used 27 out of 71 classification datasets from the standard AutoML benchmark (Gijsbers et al., 2024), which consists of curated tabular datasets from OpenML (Vanschoren et al., 2014). First, to avoid too large input prompts as well as extensive compute requirements, we selected all datasets with up to $100$ features, 100 000100000100\,000100 000 samples, and $10$ classes – resulting in $36$ datasets. We had to remove the yeast dataset due to insufficient samples per class for 10-fold cross-validation. Of the remaining $35$ available datasets, $8~{}({\sim}23\%)$ failed our memorization tests (see Appendix E) with at least one LLM, making them unsuitable for further evaluation – resulting in 27 datasets.
Evaluation Setup. For each dataset, we perform 10-fold cross-validation. For each fold, we run OpenFE and prompt each LLM to generate $20$ features. We assess the predictive performance of feature engineering following Zhang et al. (2023) by evaluating LightGBM (Ke et al., 2017) on the original features and the original features plus the newly generated features. Note, due to using a feedback loop, we only add features to the dataset when they improve the predictive performance on validation data (see Appendix A) We measured predictive performance using ROC AUC. Moreover, we mitigate a positional bias of our prompt template by repeating feature generation five times with an arbitrary order of operators. Finally, we compute the frequencies of operators across all new features, in total 27 0002700027\,00027 000.
5 Results
@@
Figure 2: Frequency of Feature Engineering Operators. We present the frequency of how often an operator was used to create a new feature across all datasets, folds, and repetitions for OpenFE, GPT-4o-mini, Gemini-1.5-flash, Llama3.1 8B, and Mistral 7B. The highest frequency observed for OpenFE is ${\sim}13.42\%$ with the complex operator GroupByThenRank. In contrast, for GPT-4o-mini, it is ${\sim}32.27\%$ with the simple operator Multiply.
@@
Table 1: Names and Frequencies of the Most and Least Frequent Operator. We present the most frequent (Max Freq.) and least frequent (Min Freq.) operator and their frequency per method/model. Each LLM model has a higher maximal and lower minimal frequency than OpenFE.
@@
Table 2: Operators Making Up $90\%$ of the Total Distribution. We show the set of operators that make up $90\%$ of the frequency distribution. OpenFE and both small open-source models require $10$ operators to obtain $90\%$ while GPT-4o-mini takes only $5$.
@@
Table 1 presents names and frequencies of the most frequently generated operators by OpenFE and each LLM. Surprisingly, the small open-source LLMs most often select a complex operator, while both big LLMs favor simple operators. Nevertheless, the frequencies for the most used operators are significantly higher for LLMs than those observed for OpenFE. We further this analysis with Table 2, which shows all operators required to accumulate $90\%$ of the total distribution. Notable, GPT-4o-mini features only five operators, with four of them - add, subtract, multiply, divide - representing basic arithmetic operators. This highlights a lack of complexity in the applied operators of one of the most complex LLMs.
@@
In some cases, LLMs could not follow the instructions of our prompt template for generating a new feature. In these cases, the LLM usually proposed an operator not on the list of allowed operators. We show the frequency of occurrences for such invalid-operators in Figure 2, represented by the right-most operator. Mistral-7B, exhibits the highest frequency of invalid-operator across all LLMs, with the frequency even exceeding other allowed operators and being part of the ten most used operators for this LLM, as shown in Table 2. While concerning, failures for code-generation-based feature engineering methods of a similar model size are still much higher; see (Hirose et al., 2024).
We analyzed the bias of LLMs across a collection of datasets to obtain a meaningful conclusion. However, practitioners likely want to know if an LLM exhibits a bias for their data. Therefore, we highlight two of the 27 datasets as an example for practitioners. Figure 3 shows the operator frequencies for the "Amazon_employee_access" dataset. We observe that the LLMs exhibit a very different distribution from OpenFE, indicating a strong bias of LLMs. While the features engineered by the LLMs match each other, only Mistral-7B engineers features similar to OpenFE. We highlighted this dataset because it was the only one where LLMs seem to favor complex features more than simple ones. This indicates that the dataset, and its information presented in the prompt, contribute to the bias exhibited by LLMs. For the second example, we show the operator frequencies for the "phoneme" dataset in Figure 4. The features engineered by the LLMs were also found to be optimal operators by OpenFE. However, GPT-4o-mini strongly prefers Add, similar to the bias observed across all datasets. Yet, the LLMs rarely use the one complex operator frequently found by OpenFE (Resiudal). This shows that an analysis across datasets is required to avoid dataset-specific noise.
@@
Figure 3: Frequency of Feature Engineering Operators for a Dataset With Strong Bias. We visualize the frequencies of operators for the "Amazon_employee_access" dataset. OpenFE favors operators to the left, while the LLMs focus on the operators to the right. For this dataset, LLMs favor complex operators (e.g., GroupByThenMean) but also frequently suggest Add and Subtract. OpenFE suggest simple operators most often but also GroupByThenRank. Mistral-7B can match the suggestions of OpenFE better than other LLMs.
@@
Figure 4: Frequency of Feature Engineering Operators for a Dataset Without Strong Bias. We visualize the frequencies of operators for the "phoneme" dataset. The distribution of the LLMs and OpenFE is reasonably well aligned, except for GPT-4o-mini for Add. OpenFE and all LLMs frequently create new features with Divide and Subtract. Nevertheless, the LLMs still fail to use the complex operator Residual, as done by OpenFE.
@@
Figure 5 and Table 3 present the predictive accuracy of the features engineered with each LLM and OpenFE. We present raw results per dataset in Appendix F. The results demonstrate the negative impact of the bias on the quality of the generated features. The two big frontier models, which are more biased toward simple operators, perform worse on average. While the two small open-source models improve performance, but still perform worse than OpenFE.
Table 3: Predictive Performance Improvement With Feature Engineering. We show the number of datasets with improvements and the average relative improvement for OpenFE and each LLM.
@@
Figure 5: Relative Improvements of Feature Engineering. We visualize the distribution of relative improvements using boxplots for OpenFE and each LLM. We compute improvement relative to a LightGBM classifier trained only on the original dataset (red horizontal line). A higher relative improvement indicates that the performance of LightGBM improved when training on the original data plus the new features generated by OpenFE or an LLM. OpenFE improves the performance on most datasets and has the highest median relative improvement, as shown by the black horizontal line in the box. In contrast, Gemini-1.5-flash rarely improves the performance.
@@
Figure 6: Frequency of Feature Engineering Operators for Random Search with LLMs. Distribution over the frequency of feature engineering operators when simulating random search by masking the operators’ names in the prompt. The frequency denotes how often an operator was used to create a new feature across all datasets, folds, and repetitions for GPT-4o-mini, Gemini-1.5-flash, Llama3.1 8B, and Mistral 7B. Compared to Figure 2, the distribution for all models exhibits a significantly higher degree of uniformity, and the previously observed bias toward simple operators does not manifest. Notably, Mistral 7B has again the highest frequency of generating invalid operators.
@@
We additionally investigate whether our prompt template influenced our results. Therefore, we adapt our prompt template to mirror a random search. That is, we force an LLM to generate new features by randomly selecting the most appropriate operator. To this end, we employed the same experimental setup as our primary experiments. However, we masked the actual names of the operators in the instructions prompt. Each operator is assigned a numeric label, which the LLM selects. Subsequently, these numeric labels are mapped back to the names of the corresponding operators. Notably, we again shuffle the order of operators five times.
Figure 6 shows the distribution over the frequency of operators with LLM-based random search for feature engineering. We observe that all models exhibit a significantly higher degree of uniformity compared to Figure 2. Yet, we do not observe total uniformity as expected for a random search, which aligns with observations for LLMs by Hopkins et al. (2023). Moreover, the previously observed bias toward simple operators does not manifest anymore. This is particularly visible for GPT-4o-mini. Mistral 7B has again the highest selection frequency of invalid operator labels, i.e., fails to follow the prompt’s instructions. We conclude that our prompt template did not cause the bias toward simple operators. Instead, the content of the prompt, in combination with the LLM, causes the bias.
6 Conclusion

### v2

@@ Abstract.
Tabular machine learning problems often require time-consuming and labor-intensive feature engineering. Recent efforts have focused on leveraging large language models’ (LLMs) domain knowledge to tackle feature engineering. At the same time, researchers have observed ethically concerning negative biases in other LLM-related use cases, such as text generation. These developments raise concerns about the generalizability and performance of LLMs in automated data science and motivate us to investigate whether LLMs exhibit a bias that adversely affect feature engineering performance. While not ethically concerning, such a bias could hinder practitioners from fully utilizing LLMs for automated data science. Therefore, we propose a novel method to systematically evaluate bias by analyzing anomalies in the frequency of feature construction operators (e.g., addition, grouping, combining) suggested by LLMs when engineering new features. Our experiments cover the bias evaluation of four LLMs, two big frontier and two small open-source models, across 27 tabular datasets. Our results indicate that LLMs are biased toward simple operators, such as addition, and can fail to utilize more complex operators, such as grouping followed by aggregations. Furthermore, the bias can negatively impact the predictive performance when using LLM-generated features. Our results highlight the need to mitigate biases in LLM-driven feature engineering and be cautious about their applications to unlock their full potential for automated data science.
LLM, Feature Engineering, Bias, Tabular Data
1. Introduction
@@
Machine learning problems involving tabular data span across many domains, such as medical diagnosis, cybersecurity, and fraud detection (Borisov et al., 2022a; van Breugel and van der Schaar, 2024). The original data for these problems (e.g., an Excel sheet) often requires manual feature engineering by a domain expert to solve the machine learning problem accurately (Tschalzev et al., 2024). During (automated) feature engineering, various operators (e.g., Add, Divide, GroupByThenMean) are applied to existing features to create new features that enhance predictive performance (Kanter and Veeramachaneni, 2015; Prado and Digiampietri, 2020; Mumuni and Mumuni, 2024).
@@
Large language models (LLMs) understand various domains (Kaddour et al., 2023; Kasneci et al., 2023; Hadi et al., 2024), tabular data (Ruan et al., 2024; Fang et al., 2024), and feature engineering (Hollmann et al., 2024; Jeong et al., 2024; Malberg et al., 2024). Leveraging their capabilities, tools such as CAAFE (Hollmann et al., 2024) or OCTree (Nam et al., 2024) have enabled practitioners to automate the labor intensive process of feature engineering with the help of LLMs, significantly reducing the time and effort required. Thus, data scientists have started to leverage LLMs for feature engineering, especially via the use of CAAFE (Hollmann et al., 2024), a powerful method for automatic feature engineering with LLMs. [^1] For example, see these recent Kaggle competition write-ups (Hatch, 2024; Türkmen, 2024).
@@
Despite their utility, LLMs are known to have negative biases as observed for chat applications (Kotek et al., 2023; Navigli et al., 2023; Gallegos et al., 2024; Bang et al., 2024) or when ”meticulously delving” into text generation (Liang et al., 2024). In the context of automated data science, these biases, may also hinder the performance of LLMs in automated data science workflows. Motivated by this observation, we investigate whether LLMs exhibit a bias that negatively impacts the quality of their engineered features. If a bias is found and can be circumvented, LLMs would become a more potent tool for data scientists.
@@
To identify whether a bias exists, we inspect the frequency of operators LLMs use when engineering features. This parallels inspecting the frequency of words to detect LLM-generated text (Liang et al., 2024). Assuming LLMs use their world knowledge and reasoning capabilities to employ the most appropriate operator, we expect the operators’ frequencies to be similar to those of the optimal operators. For example, if adding two features is often the optimal feature, an LLM should frequently add two features. Moreover, if the LLM does not know the optimal operator, it should resort to a random search over operators.
@@
Therefore, we introduce a method to compare the frequency distribution of feature engineering operators used by an LLM to the frequencies of the same operators obtained by searching for the optimal features with automatic black-box feature engineering using OpenFE  (Zhang et al., 2023). Using this method, we evaluated the bias of 4 LLMs, namely GPT-4o-mini (OpenAI, 2024), Gemini-1.5-flash (Gemini-Team, 2024), Llama3.1-8B (Touvron et al., 2023), and Mistral7B-v0.3 (Jiang et al., 2023). In an additional experiment, we also evaluated the bias of the more powerful model GPT-4o (OpenAI, 2024). We obtained the distribution over operators for 27 tabular classification datasets unknown to all LLMs.
@@
Our results demonstrate that LLMs can have a negative bias when engineering features for tabular data. LLMs favor simple operators during feature engineering (e.g., Add) and rarely use more complex operators (e.g., GroupByThenMean). In contrast, automatic black-box feature engineering favors complex operators but also uses simple operators.
@@
In particular, we observed a strong bias and negative impact for GPT-4o-mini and Gemini-1.5-flash, two big frontier models (Chiang et al., 2024). Both models select simple operators most often and their generated features decrease the average predictive performance. In contrast, Llama3.1-8B and Mistral7B-v0.3, the small open-source models, are less biased or negatively impacted. Nevertheless, none of the LLMs generates features following distribution over operators obtained by OpenFE. Likewise, the features generated by OpenFE improve the predictive performance on average the most.
@@
2. Related Work
Feature Engineering Without Large Language Models. Previous work dedicated considerable effort toward automating the process of feature engineering (Kanter and Veeramachaneni, 2015; Prado and Digiampietri, 2020; Mumuni and Mumuni, 2024). Various black-box methods have been proposed, such as ExploreKit (Katz et al., 2016), AutoFeat (Horn et al., 2020), BioAutoML (Bonidia et al., 2022), FETCH (Li et al., 2023), and OpenFE (Zhang et al., 2023). These methods typically generate new features in two steps: 1) create a large set of candidate features by applying mathematical (e.g., Add) or functional (e.g., GroupByThenMean) operators to features, and 2) return a small set of promising features selected from all candidate features.
Feature Engineering with Large Language Models. LLMs allow us to exploit their (potential) domain knowledge for feature engineering and can act as a proxy to a domain expert or data scientist during the feature engineering process. To illustrate, LLMs can be prompted to suggest code for generating new features (Hollmann et al., 2024; Hirose et al., 2024), to select predictive features (Jeong et al., 2024), or to use rule-based reasoning for generation (Nam et al., 2024). Most notably, CAAFE (Hollmann et al., 2024) presents a simple yet effective method to generate new features by proposing Python code to transform existing features in the dataset into new valuable features. This method lays the foundation to our proposed feature generation method, thanks to its wide application in practice [^2] For example, see these recent Kaggle competition write-ups (Hatch, 2024; Türkmen, 2024). and also in research (Malberg et al., 2024; Guo et al., 2024; Zhang et al., 2024a).
@@ 2. Related Work
Other Applications of Large Language Models for Tabular Data. Many researchers recently started using LLMs for applications related to tabular data. To avoid confusion in this plethora of recent work, we highlight similar but not directly related work to our contribution. Our work is not directly related to LLMs on tabular question answering (Ghosh et al., 2024; Grijalba et al., 2024; Wu et al., 2024), tabular dataset generation (Borisov et al., 2022b; van Breugel et al., 2024; Panagiotou et al., 2024), tabular data manipulation (Zhang et al., 2024b; Qian et al., 2024; Lu et al., 2024), or tabular few-shot predictions (Hegselmann et al., 2023; Han et al., 2024; Gardner et al., 2024)
3. Method: Analyzing Feature Engineering Bias
Figure 1. Our Method to Analyze Feature Engineering Bias of an LLM. Our method is split into three stages. In the first stage A), we discard tabular datasets, which the LLM might have memorized. In the second stage B), we instruct the LLM to select the best operators to engineer new features. At the same stage, we use OpenFE (Zhang et al., 2023), black-box automated feature engineering, to determine a proxy for the optimal operators. In the third stage C), we compare the frequencies of operators for the LLM and OpenFE. That is, we look for anomalies when contrasting the distributions, such as using simple operators much more than complex ones.
@@ 3. Method: Analyzing Feature Engineering Bias
The LLM is then instructed to generate precisely one new feature by selecting one or two existing features and applying one of the available operators. We employ chain-of-thought (CoT) prompting (Wei et al., 2023) to boost the expressive power of the LLM (Li et al., 2024). In addition, the LLM explains why a feature was generated with CoT. We also employ a feedback loop, to pre-evaluate proposed features, similar to CAAFE (Hollmann et al., 2024), which we detail in Appendix A.
@@
At its core, our objective is to compare the distribution of operators proposed by the LLM with the distribution of the optimal operators. That said, we do not know the optimal operators for a dataset. Thus, as a proxy, we use the operators suggested by OpenFE (Zhang et al., 2023).
@@ 4. Experiments
Large Language Models. We use four LLMs hosted by external providers via APIs. In detail, we used GPT-4o-mini (OpenAI, 2024) and Gemini-1.5-flash (Gemini-Team, 2024) to represent big frontier LLMs and Llama3.1 8B (Touvron et al., 2023) and Mistral 7B Instruct v0.3 (Jiang et al., 2023), hosted by Together AI [^4] https://www.together.ai/, to represent small open-source models. The API usage cost ${\sim}200\$$.
@@
Datasets. We used 27 out of 71 classification datasets from the standard AutoML benchmark (Gijsbers et al., 2024), which consists of curated tabular datasets from OpenML (Vanschoren et al., 2014). In order to avoid too large input prompts as well as extensive compute requirements, we selected all datasets with up to $100$ features, 100 000100000100\,000100 000 samples, and $10$ classes – resulting in $36$ datasets. We had to remove the yeast dataset due to insufficient samples per class for 10-fold cross-validation. Of the remaining $35$ available datasets, $8~{}({\sim}23\%)$ failed our memorization tests (see Appendix E) with at least one LLM, making them unsuitable for further evaluation – resulting in 27 datasets.
Evaluation Setup. For each dataset, we perform 10-fold cross-validation. For each fold, we run OpenFE and prompt each LLM to generate $20$ features. We assess the predictive performance of feature engineering following Zhang et al. (2023) by evaluating LightGBM (Ke et al., 2017) on the original features and the original features plus the newly generated features. Note, that due to using a feedback loop, we only add features to the dataset when they improve the predictive performance on validation data (see Appendix A) We measured predictive performance using ROC AUC, as it relies on prediction probabilities rather than direct labels (Provost et al., 1998).
Furthermore, we mitigate a positional bias of our prompt template by repeating feature generation five times with an arbitrary order of operators. Finally, we compute the frequencies of operators across all new features, in total 27 0002700027\,00027 000.
5. Results
@@
Figure 2. Frequency of Feature Engineering Operators. We present the frequency of how often an operator was used to create a new feature across all datasets, folds, and repetitions for OpenFE, GPT-4o-mini, Gemini-1.5-flash, Llama3.1 8B, and Mistral 7B. The highest frequency observed for OpenFE is ${\sim}13.42\%$ with the complex operator GroupByThenRank. In contrast, for GPT-4o-mini, it is ${\sim}32.27\%$ with the simple operator Multiply.
@@
Table 1. Names and Frequencies of the Most and Least Frequent Operator. We present the most frequent (Max Freq.) and least frequent (Min Freq.) operator and their frequency per method/model. Each LLM model has a higher maximal and lower minimal frequency than OpenFE.
@@
Table 2. Operators Making Up $90\%$ of the Total Distribution. We show the set of operators that make up $90\%$ of the frequency distribution. OpenFE and both small open-source models require $10$ operators to obtain $90\%$ while GPT-4o-mini takes only $5$.
@@
Table 1 presents names and frequencies of the most frequently generated operators by OpenFE and each LLM. Surprisingly, the small open-source LLMs most often select a complex operator, while both big LLMs favor simple operators. Nevertheless, the frequencies for the most used operators are significantly higher for LLMs than those observed for OpenFE.
We further this analysis with Table 2, which shows all operators required to accumulate $90\%$ of the total distribution. Notable, GPT-4o-mini features only five operators, with four of them - add, subtract, multiply, divide - representing basic arithmetic operators. This highlights a lack of complexity in the applied operators of one of the most complex LLMs.
In some cases, LLMs could not follow the instructions of our prompt template for generating a new feature. In these cases, the LLM usually proposed an operator not on the list of allowed operators. We show the frequency of occurrences for such
invalid-operators in Figure 2, represented by the right-most operator. Mistral-7B, exhibits the highest frequency of
@@
invalid-operator across all LLMs, with the frequency even exceeding other allowed operators and being part of the ten most used operators for this LLM, as shown in Table 2. While concerning, failures for code-generation-based feature engineering methods of a similar model size are still much higher; see (Hirose et al., 2024).
@@
Figure 3. Frequency of Feature Engineering Operators for a Dataset With Strong Bias. We visualize the frequencies of operators for the ”Amazon_employee_access” dataset. OpenFE favors operators to the left, while the LLMs focus on the operators to the right. For this dataset, LLMs favor complex operators (e.g., GroupByThenMean) but also frequently suggest Add and Subtract. OpenFE suggest simple operators most often but also GroupByThenRank. Mistral-7B can match the suggestions of OpenFE better than other LLMs.
We analyzed the bias of LLMs across a collection of datasets to obtain a meaningful conclusion. However, practitioners likely want to know if an LLM exhibits a bias for their data. Therefore, we highlight two of the 27 datasets as an example for practitioners. Figure 3 shows the operator frequencies for the ”Amazon_employee_access” dataset. We observe that the LLMs exhibit a very different distribution from OpenFE, indicating a strong bias of LLMs. While the features engineered by the LLMs match each other, only Mistral-7B engineers features similar to OpenFE. We highlighted this dataset because it was the only one where LLMs seem to favor complex features more than simple ones. This indicates that the dataset, and its information presented in the prompt, contribute to the bias exhibited by LLMs.
@@
Figure 4. Frequency of Feature Engineering Operators for a Dataset Without Strong Bias. We visualize the frequencies of operators for the ”phoneme” dataset. The distribution of the LLMs and OpenFE is reasonably well aligned, except for GPT-4o-mini for Add. OpenFE and all LLMs frequently create new features with Divide and Subtract. Nevertheless, the LLMs still fail to use the complex operator Residual, as done by OpenFE.
For the second example, we show the operator frequencies for the ”phoneme” dataset in Figure 4. The features engineered by the LLMs were also found to be optimal operators by OpenFE. However, GPT-4o-mini strongly prefers Add, similar to the bias observed across all datasets. Yet, the LLMs rarely use the one complex operator frequently found by OpenFE (Residual). This shows that an analysis across datasets is required to avoid dataset-specific noise.
Figure 5. Relative Improvements of Feature Engineering. We visualize the distribution of relative improvements using boxplots for OpenFE and each LLM. We compute improvement relative to a LightGBM classifier trained only on the original dataset (red horizontal line). A higher relative improvement indicates that the performance of LightGBM improved when training on the original data plus the new features generated by OpenFE or an LLM. OpenFE improves the performance on most datasets and has the highest median relative improvement, as shown by the black horizontal line in the box. In contrast, Gemini-1.5-flash rarely improves the performance.
@@
Figure 5 and Table 3 present the relative improvements in predictive accuracy of the features engineered with each LLM and OpenFE in comparison to a system without feature engineering. We present raw results per dataset in Appendix F. Furthermore we evaluate the statistical significance of these results in Appendix G. The results demonstrate the negative impact of the bias on the quality of the generated features. The two big frontier models, which are more biased toward simple operators, perform worse on average. While the two small open-source models improve performance, but still perform worse than OpenFE. Additionally, the effectiveness of OpenFE is highlighted, improving predictive accuracy on 21 of 27 benchmark datasets.
Table 3. Predictive Performance Improvement With Feature Engineering. We show the number of datasets with improvements and the average relative improvement for OpenFE and each LLM.
@@
Figure 6. Frequency of Feature Engineering Operators for Random Search with LLMs. Distribution over the frequency of feature engineering operators when simulating random search by masking the operators’ names in the prompt. The frequency denotes how often an operator was used to create a new feature across all datasets, folds, and repetitions for GPT-4o-mini, Gemini-1.5-flash, Llama3.1 8B, and Mistral 7B. Compared to Figure 2, the distribution for all models exhibits a significantly higher degree of uniformity, and the previously observed bias toward simple operators does not manifest. Notably, Mistral 7B has again the highest frequency of generating invalid operators.
@@
We additionally investigate whether the observed bias primarily arises from positional preferences related to the positioning of operator names in our prompt. We additionally investigate whether our prompt template influenced our results. Therefore, we adapt our prompt template to mirror a random search. That is, we force an LLM to generate new features by randomly selecting the most appropriate operator. To this end, we employed the same experimental setup as our primary experiments. However, we masked the actual names of the operators in the instructions prompt. Each operator is assigned a numeric label, which the LLM selects. Subsequently, these numeric labels are mapped back to the names of the corresponding operators. Notably, we again shuffle the order of operators five times.
Figure 6 shows the distribution over the frequency of operators with LLM-based random search for feature engineering. We observe that all models exhibit a significantly higher degree of uniformity compared to Figure 2. Yet, we do not observe total uniformity as expected for a random search, which aligns with observations for LLMs by Hopkins et al. (2023). Moreover, the previously observed bias toward simple operators does not manifest anymore. This is particularly visible for GPT-4o-mini. Mistral 7B has again the highest selection frequency of invalid operator labels, i.e., fails to follow the prompt’s instructions. We conclude that the positioning the operator names in our prompt template did not cause the bias toward simple operators. Instead, the content of the prompt, in combination with the LLM, causes the bias.
Figure 7. Comparison of Operator Distributions for GPT-4o and GPT-4o-mini. We present the distributions of selected operators on 4 different benchmark datasets. Generally the distribution for GPT-4o is smoother in comparison to GPT-4o-mini. However, the problem of relying heavily on only few (simple) operators, is similar to GPT-4o-mini, as visible from the set of selected operators as well as the frequencies with which the simple operators are selected accross all 4 datases.
ADDITIONAL EXPERIMENT. USING MORE POWERFUL LLMS AS FOUNDATION.
Furthermore, we evaluate whether the apparent bias can be circumvented by selecting more powerful models as foundation. We conducted the same experiments described in Section 4 on a subset of four benchmark datasets in OpenAI’s GPT-4o model. This is due to high cost of using GPT-4o for large amount of tokens. We compare the distribution of selected operators by the GPT-4o model to the distribution of the GPT-4o-mini model, presented in Figure 7. When considering the churn and phoneme dataset, a strong similarity between the distributions of GPT-4o and GPT-4o-mini is apparent. For the ada and shuttle dataset the distribution of GPT-4o is visibly smoother in comparison to the distribution of GPT-4o-mini in regards to the frequencies with which operators are selected. However when considering the types of selected features, the bias towards simple operators is still very strongly represented by GPT-4o, indicating that the usage of more powerful models does not circumvent the bias towards simpler operators.
6. Conclusion